Circular No. 6783

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ef IN UGC 4107 J. Y. Hu, Y. L. Qiu, Q. Y. Qiao, and J. Y. Wei, Beijing Astronomi-

cal Observatory (BAO), report that the possible supernova in UGC 4107 reported on IAUC 6778 is being monitoring photometrically and spectroscopically at BAO. The object did not appear on CCD images (limiting mag 18.5) taken on Nov. 16. CCD images taken with the 0.60-m telescope show that it brightened from Nov. 28 to Dec. 8; available CCD magnitudes: Nov. 28.70 UT, 16.8; Dec. 7.74, 16.3; 8.72, 16.2. The low-dispersion spectra obtained with 2.16-m telescope on Nov. 28, Dec. 4, and 7, show that the spectra still show the features reported by Garnavich et al. (IAUC 6778), but the peaks are shifted redward. The peak at 510 nm reported by Garnavich et al. shifted to 520, 535, and 545 nm in spectra obtained on Nov. 28, Dec. 4, and 7, respectively. The Na D absorption lines in SN 1997ef show the same redshift as does the Hα from UGC 4107, implying that the object is located in the host galaxy.

A. V. Filippenko, University of California at Berkeley (UCB), reports that he has examined a spectrum (range 470–1020 nm, resolution 1.5 nm) of SN 1997ef, obtained by E. L. Martín (UCB) on Dec. 10 with the 3-m Shane reflector at Lick Observatory: “The spectrum superficially resembles that of a supernova but in detail is unlike that of any known type of supernova; the features are unusually broad (FWHM 50 nm). The strongest emission line is at rest wavelength 540 nm; if identified with [O I] 557.7 nm, its blueshift is presumably due to obscuration of the object’s back side. A smaller peak at 639 nm is then consistent with Hα blueshifted by the same amount. Weak peaks are visible at 777 nm (possibly unshifted O I 777.4 nm), 850 nm (slightly shifted Ca II?), and 909 nm (shifted O I 926 nm?). All of the maxima listed by Garnavich et al. (IAUC 6778) have moved substantially to the red. If this offset is applied to their reported peak at 440 nm, the feature would now be roughly consistent with unshifted Mg I] 457.1 nm. Aside from the tentative Hα identification (which may instead be unshifted [O I]), the spectrum vaguely resembles what might be expected from a denuded O-Ne-Mg (or perhaps C-O) core. Thus, the object could be a previously unobserved, extreme example of a stripped (type-Ic-like) supernova, though other interpretations are possible. The great breadth of the lines suggests a low mass for the ejecta. Further observations of this unique object are encouraged.”

Unfiltered CCD magnitude estimates: Nov. 8 UT, [17.5: (W. D. Li, R. R. Treffers, Filippenko, and M. Modjaz, Lick Observatory); Dec. 2.4, 16.5: (Li et al.); 7.98, 16.7 (S. Moretti and S. Tomaselli, Forlí, Italy).

1997 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green